OpenStudy (anonymous):

will give metal to best answer. order 3.21 x 10^6, 42 x 10^4, and 0.11 x 10^10 from least to greatest

OpenStudy (mathstudent55):

First, write all numbers in scientific notation. Then compare them.

OpenStudy (rane):

they are already in scientific notation

OpenStudy (rane):

0.11*10^10, 3.21*10^6, 42*10^4

OpenStudy (mathstudent55):

Not really. A number in scientific notation consists of a number from 1 to 10, not including 10, multiplied by an integer power of 10. 3.21 x 10^6 This number is in scientific notation. 42 x 10^4 This number is not in scientific notation because 42 is not between 1 and 10. 0.11 x 10^10 This is also not in scientific notation because 0.11 is not between 1 and 10. Let's write all numbers in scientific notation: 3.21 x 10^6 = 3.21 x 10^6 42 x 10^4 = 4.2 x 10^5 0.11 x 10^10 = 1.1 x 10^9 Now we can put them in order by comparing the powers of 10. From smallest to largest, they are: 42 x 10^4, 3.21 x 10^6, 0.11 x 10^10

OpenStudy (rane):

thank you, u r right

OpenStudy (mathstudent55):

wlcm